three identical spin half fermions are to be distributed in three non degenerate distinct energy level the no of ways this can be done is = ?
1
Expert's answer
2020-01-24T05:18:07-0500

n=(2Nl)!/(Nl!Nf!)=(2×3)!/(3!×3!)=6!/(6×6)=720/36=20.

Answer: 20.